A project costs 29,000

and is expected to have a useful life of three years of which its scrap value will be 5000.

The project is expected to yield net profits of 1,000p.a. over its useful life.

Using the average book value of the asset,

the accounting rate of return (ARR) will be____?


1
Expert's answer
2021-12-19T18:09:16-0500

Initial cost = 29,000

Scrap value = 5000

Useful life =3 years

"Depreciation=\frac{Initial cost-scrap value}{useful life}"

"Depreciation=\frac{29000-5000}{3}=8000"

"ARR=\frac{1000-8000}{29000}=-24.14\%"


The accounting rate of return (ARR) will be -24.14%.